OpenStudy (aravindg):

Few questions I have doubts on:

OpenStudy (aravindg):

1. Two simple pendulums of length 1 m and 4m are both given small displacement in same direction at the same instant. They will be again in phase after the short pendulum has completed number of oscillations=.....

OpenStudy (aravindg):

\[2T_2=T_1\] \[2 \pi f_1=2(2 \pi f_2)\]

OpenStudy (aravindg):

Now what to do?

OpenStudy (vincent-lyon.fr):

2 periods for the smaller one = 1 period for the longer one so the answer is 2.

OpenStudy (aravindg):

Oh is that all? I was confused.

OpenStudy (vincent-lyon.fr):

Unless what they call "oscillation" is not a full period but half a period. In that case, the answer would be 4.

OpenStudy (aravindg):

A bullet of mass 10 g and speed 500 m/s is fired into a door and gets embedded at exactly center of the door. The door is 1 m wide and weighs 12kg. It is hinged at one end and rotates in a vertical axis. The angular speed of door just after bullet embeds it will be?

OpenStudy (vincent-lyon.fr):

Use conservation of angular momentum. 0.010kg x 500 x 0.5m = I x omega with I = m L²/3 = 12 x 1² /3 = 4 kg.m² omega = 0.6 rad/s

OpenStudy (aravindg):

So for the bullet door system angular momentum is conserved always?

OpenStudy (vincent-lyon.fr):

Yes because there is no net torque acting and the collision time is infinitesimal.

OpenStudy (aravindg):

Orbits of a particle moving in a circle are such that perimeter of orbit=nlambda where lambda is de broglie wavelength. For a charged particle moving in a plane perpendicular to mag field radius of nth orbit will be proportional to?

OpenStudy (aravindg):

Ok I know how to do this qn. my doubt is why the general eqn \[r \alpha n^2 \] Does not hold here.

OpenStudy (vincent-lyon.fr):

I do not understand. Is it alpha between r and n² ?

OpenStudy (aravindg):

I mean r proportional to n^2 for an electron moving around nucleus.

OpenStudy (vincent-lyon.fr):

It must be different here, as the particle is not orbiting an atom with an 1/r² force but is moving in a magnetic field. I'm not sure. I'll think about it. See you.

OpenStudy (aravindg):

ok bye. Thanks again!